# Exploring the Adversarial Frontier: Quantifying Robustness via Adversarial Hypervolume

### v1

@@ Abstract
The escalating threat of adversarial attacks on deep learning models, particularly in security-critical fields, has underscored the need for robust deep learning systems. Conventional robustness evaluations have relied on adversarial accuracy, which measures a model’s performance under a specific perturbation intensity. However, this singular metric does not fully encapsulate the overall resilience of a model against varying degrees of perturbation. To address this gap, we propose a new metric termed adversarial hypervolume, assessing the robustness of deep learning models comprehensively over a range of perturbation intensities from a multi-objective optimization standpoint. This metric allows for an in-depth comparison of defense mechanisms and recognizes the trivial improvements in robustness afforded by less potent defensive strategies. Additionally, we adopt a novel training algorithm that enhances adversarial robustness uniformly across various perturbation intensities, in contrast to methods narrowly focused on optimizing adversarial accuracy. Our extensive empirical studies validate the effectiveness of the adversarial hypervolume metric, demonstrating its ability to reveal subtle differences in robustness that adversarial accuracy overlooks. This research contributes a new measure of robustness and establishes a standard for assessing and benchmarking the resilience of current and future defensive models against adversarial threats.
@@ I Introduction
The increasing emphasis on evaluating the inherent robustness of deep learning (DL) models and developing more robust DL systems constitutes a pivotal area of research within the machine learning (ML) community [1, 2, 3, 4, 5]. This paradigm shift is primarily driven by growing security concerns associated with adversarial attacks on deep neural network-based (DNN) image classification systems. These attacks introduce imperceptible perturbations to input images, which are typically indistinguishable from the human eye, leading to incorrect classifications by the DNN model [6, 7, 8]. The potential implications of such vulnerabilities are particularly alarming in contexts critical to security, such as face recognition [9] and autonomous driving [10], where the consequences of misclassification could be severe. Nonetheless, efforts to construct adversarially robust models face substantial challenges, particularly the absence of a comprehensive evaluation framework for rigorously evaluating the robustness of DL models.
@@
Conventionally, adversarial accuracy has been the principal metric for robustness evaluation, measuring a model’s resilience to adversarial examples generated through a range of attack strategies [7, 11, 12]. The emergence of increasingly sophisticated attacks, such as the Fast Gradient Sign Method (FGSM) [7], DeepFool [13], Projected Gradient Descent (PGD) [1], and Auto Projected Gradient Descent (APGD) [14], has led to an expansion of benchmark adversarial accuracies. Despite this, the field lacks a unified robustness metric due to the differing adversarial accuracy results across various attacks. Notably, the integration of APGD [14] and Square Attack [15] within AutoAttack [14] has established a widely recognized benchmark for adversarial accuracy assessment. The development of standard toolkits like RobustBench [16] has also substantially advanced the uniform comparison of adversarial defense capabilities across different defensive models, positioning the adversarial accuracy of the models under AutoAttack as the gold standard.
@@
Recent research has challenged the adequacy of adversarial accuracy as the robustness metric for DL models, demonstrating that it fails to capture the full scope of a model’s resilience against attacks [11, 17, 18]. Robey et al. [17] have proposed the probability of misclassification as a more comprehensive measure of robustness, calculated through random sampling, although they acknowledge the computational challenges posed by the complexity of high-dimensional spaces. Oliver et al. [18] have explored non-adversarial spaces using geometric methods, assessing the extent of these regions within the sample space as an indicator of robustness. Despite these contributions, one can argue that neither misclassification probability nor adversarial region size sufficiently represent a model’s robustness, as they fail to consider the variation in a model’s response to varying levels of perturbation intensities.
@@
To address the shortcomings of existing robustness metrics, which focus solely on model robustness at a fixed perturbation level, we introduce a new metric called adversarial hypervolume. This metric evaluates DL model robustness by examining the normalized confidence scores of worst-case adversarial examples across various levels of perturbation intensities. It facilitates meaningful comparisons between defensive models with comparable adversarial accuracy and evaluation of the incremental robustness improvements provided by weaker defensive measures like Feature Squeezing and JPEG compression [19]. A higher adversarial hypervolume metric value indicates improved model robustness throughout various perturbation intensities, thereby offering a more comprehensive assessment of a model’s resilience against adversarial attacks. We illustrate a conceptual comparison between our proposed metric and two existing metrics in Figure 1.
@@
Our metric, adversarial hypervolume, is based on the concept of hypervolume [20] from the multi-objective optimization literature [21, 22, 23]. We begin by constructing a multi-objective optimization problem that mirrors the complexity of actual attack scenarios, with the dual aims of minimizing the perturbation’s magnitude and simultaneously reducing the model’s confidence in the original classification. In contrast, much of existing research approaches the problem from the perspective of single-objective optimization with a distance constraint [7, 1] or employs a weighted sum of multiple objectives [11]. We subsequently delineate a set of adversarial examples that illustrate the model’s unique vulnerability landscape when subjected to an attack. Each example signifies a nadir of confidence corresponding to a particular magnitude of perturbation, thereby delineating the adversarial frontier. This concept is inspired by the Pareto optimality principle [21]. Ultimately, we quantify a model’s robustness by calculating the adversarial hypervolume, which measures the size of the region demarcated by the adversarial frontier.
@@
  • We introduce a novel metric, the adversarial hypervolume, which quantifies the robustness of DL models against a range of perturbation intensities, thus offering a comprehensive evaluation of a model’s resilience to adversarial attacks.
  • Employing this metric, we propose a straightforward method for its computation and adopt an efficient training algorithm to enhance the robustness of DL models.
  • Our empirical findings reveal that the adversarial hypervolume effectively measures the enhancement of robustness from weaker defensive mechanisms (i.e., input transformations) and delivers important supplementary insights, such as average confidence variation, which serve as valuable adjuncts to adversarial accuracy.
  • By conducting comprehensive experiments, we have assessed the robustness of cutting-edge state-of-the-art (SOTA) defensive models, thereby setting a benchmark for future studies.
@@ II-A Adversarial Attacks and Defenses
Adversarial Attacks. Since the discovery of adversarial vulnerabilities in DL models [6], a substantial body of research has focused on the development of attack algorithms within various threat models to critically evaluate the robustness of these models [24, 25, 26]. Specifically, researchers have developed white-box attacks by leveraging complete access to a model’s parameters and gradients; in contrast, black-box attacks are devised with merely an understanding of the model’s input-output interface available to the attacker.
White-box Attacks. Exploiting full knowledge of the model, including its gradients, white-box attacks evaluate a model’s robustness by formulating and solving an optimization problem [1]. This practice has led to the development of a series of attacks, including the PGD attack [1], the Carlini-Wagner (CW) attack [11], the FGSM attack [7], the Basic Iterative Method (BIM) attack [8], and the DeepFool attack [13]. A model’s adversarial accuracy, when subjected to these attacks, is a critical metric for determining its robustness.
Black-box Attacks. Despite the efficacy of white-box attacks in assessing robustness, their real-world applicability is often restricted due to limited access to internal model details. Therefore, many black-box attacks have been developed to overcome this limitation [27, 28, 29, 30, 31, 32]. Among them, the Square attack stands out, successfully breaching the SOTA defenses using a gradient-free strategy and even surpassing some gradient-based attacks that have access to comprehensive information [14]. Within the AutoAttack framework [14], it is a critical component, utilized alongside the APGD attack to measure a model’s adversarial accuracy.
@@
Adversarial Defenses. Defensive strategies against adversarial attacks are designed to enhance a model’s inherent robustness or alleviate the effects of adversarial attacks. To improve the robustness, a range of adversarial training techniques have been introduced [1, 33]. Furthermore, several tactics have been devised to detect or neutralize adversarial examples, including input transformation [34, 19], adversarial purification [35, 36], and stateful defenses [37, 38]. Given that adversarial purification and stateful defenses are tailored for real-world classification systems rather than a standalone model, our study concentrates primarily on the robustness improvement brought by adversarial training and input transformation techniques.
@@
Adversarial Training. Adversarial training represents a prevalent approach to increase the robustness of DL models by incorporating adversarial examples into the training dataset, which enables the model to learn more robust decision boundaries [1]. Adversarial examples are often generated using the PGD attack [1], and a significant volume of research corroborates its efficacy [39, 40, 41].
@@
Input transformation. Input transformation is a cost-effective defensive strategy that can be seamlessly incorporated into systems with adversarially trained models to enhance robustness [34, 19]. However, the efficacy of input transformation has come into question, as it has been proven to be vulnerable to circumvention [12]. Moreover, using adversarial accuracy to measure the robustness improvement from input transformation techniques may lead to misleading results, since the adversarial accuracy of disparate techniques tends to be low and indistinguishable. In contrast, our research introduces the adversarial hypervolume metric, which facilitates a comparative analysis of robustness improvements by providing more detailed insights into average confidence values.
II-B Alternative Metrics for Evaluating Robustness
Although adversarial accuracy is commonly used as a primary metric to assess model robustness, researchers have proposed several alternative measures to offer a more comprehensive understanding of model robustness. These include the minimum adversarial perturbation [11], probabilistic robustness [17], and adversarial sparsity [18].
Minimal Perturbation. Measuring the smallest perturbation magnitude required to alter an input into an adversarial example is a prevalent method for evaluating the robustness of a model. Carlini et al. addressed this through an optimization problem that balances the perturbation size and the confidence of the adversarial example by considering their linear combination [11]. However, this metric fails to fully capture the spectrum of a model’s robustness (confidence score) across various perturbation levels.
Probabilistic robustness. Robey et al. introduced a metric to assess probabilistic robustness, shifting the focus from the conventional worst-case scenario [17]. Rather than just considering the adversarial accuracy, this metric evaluates the likelihood of a model’s correct predictions in the presence of adversarial perturbations. The downside is its computational intensity, often necessitating Monte Carlo simulations to approximate the adversarial example distribution, especially in high-dimensional spaces.
@@ II-B Alternative Metrics for Evaluating Robustness
Adversarial Sparsity. Recent research by Oliver et al. has presented the concept of adversarial sparsity, which probes the boundary of non-adversarial regions to determine their proportion within the entire sample space [18]. Although the method aligns with the principles of probabilistic robustness, Oliver et al. suggest a geometric method to determine the adversarial sparsity, which yields improved convergence results. Despite these advancements, measuring this metric remains challenging in high-dimensional spaces.
@@ III Adversarial Hypervolume
This section introduces the concept of adversarial hypervolume, a novel metric developed to assess the robustness of DL models against adversarial attacks. Initially, we outline the multi-objective optimization problem foundational to the adversarial hypervolume concept in Section III-A. Subsequently, we describe the adversarial frontier in Section III-B and detail the methodology for computing the adversarial hypervolume in Section III-C. We then explore the relationship between adversarial hypervolume, adversarial accuracy, and other established metrics in Section III-D.
@@ III-A Multiobjecive Adversarial Problem
Consider a classifier $f:\mathcal{X}\to\mathcal{Y}$, where $\mathcal{X}\subseteq[0,1]^{d}$ represents the input image space, and $\mathcal{Y}\subseteq[0,1]^{m}$ corresponds to the output space implicating a probability distribution across $m$ class labels. The aim of the adversarial attack is to determine the perturbation $\delta$ for a given unperturbed data point $(x,y)$ as expressed in:
$\min_{\delta}\mathcal{L}(f,x,y,\delta)=(\mathcal{L}_{\text{conf}}(f,x,y,\delta),\mathcal{L}_{\text{dist}}(\delta)).$ (1)
In this formula, $\mathcal{L}_{\text{conf}}$ denotes the confidence loss, while $\mathcal{L}_{\text{dist}}$ indicates the distance loss. Cutting-edge attacks typically utilize the negative cross-entropy loss to measure confidence loss and the $\ell_{p}$ norm distance to quantify distance loss. These attacks typically focus on minimizing an aggregated loss function or optimizing the confidence loss subject to a constraint on the distance loss [1, 11].
Motivation. The conventional approach for creating adversarial examples typically limits the secondary objective (e.g., $\|\delta\|\leq\epsilon$), focusing exclusively on optimizing the primary objective [1, 11]. However, this strategy does not thoroughly explore the loss landscape, thus failing to comprehensively evaluate the model’s robustness. Our study introduces the use of the hypervolume indicator as a comprehensive metric to assess the robustness of the model by measuring the set of optimal trade-offs between the two objectives.

### v2

@@ Abstract
The escalating threat of adversarial attacks on deep learning models, particularly in security-critical fields, has highlighted the need for robust deep learning systems. Conventional evaluation methods of their robustness have relied on adversarial accuracy, which measures the model performance under a specific perturbation intensity. However, this singular metric does not fully encapsulate the overall resilience of a model against varying degrees of perturbation. To address this issue, we propose a new metric termed as the adversarial hypervolume for assessing the robustness of deep learning models comprehensively over a range of perturbation intensities from a multi-objective optimization standpoint. This metric allows for an in-depth comparison of defense mechanisms and recognizes the trivial improvements in robustness brought by less potent defensive strategies. We adopt a novel training algorithm to enhance adversarial robustness uniformly across various perturbation intensities, instead of only optimizing adversarial accuracy. Our experiments validate the effectiveness of the adversarial hypervolume metric in robustness evaluation, demonstrating its ability to reveal subtle differences in robustness that adversarial accuracy overlooks.
@@ I Introduction
The increasing emphasis on evaluating the inherent robustness of deep learning (DL) models and developing robust DL systems constitutes a pivotal area of research within the machine learning (ML) community [1, 2, 3, 4, 5]. This trend is primarily driven by growing security concerns with adversarial attacks on deep neural network-based (DNN) image classification systems. These attacks introduce imperceptible perturbations to input images, which are typically indistinguishable from the human eye, leading to incorrect classifications by the DNN model [6, 7, 8]. Such vulnerabilities are particularly alarming in contexts critical to security, such as face recognition [9] and autonomous driving [10], where the consequences of misclassification could be severe. Nonetheless, efforts to construct adversarially robust models face substantial challenges, particularly the absence of a comprehensive evaluation metric or framework for evaluation of models’ robustness.
@@
Conventionally, adversarial accuracy has been used as the principal metric for robustness evaluation, measuring a model’s resilience to adversarial examples generated through a range of attack strategies [7, 11, 12]. The emergence of increasingly sophisticated attacks, such as the Fast Gradient Sign Method (FGSM) [7], DeepFool [13], Projected Gradient Descent (PGD) [1], and Auto Projected Gradient Descent (APGD) [14], has led to an expansion of benchmark of adversarial accuracies. Despite this, the field lacks a unified robustness metric due to the differing adversarial accuracy results across various attacks. Notably, the integration of APGD [14] and Square Attack [15] within AutoAttack [14] has established a widely recognized benchmark for adversarial accuracy assessment. The development of standard toolkits like RobustBench [16] has also substantially advanced the uniform comparison of adversarial defense capabilities across different defensive models, positioning the adversarial accuracy of the models under AutoAttack as the gold standard.
@@
However, recent research has challenged the adequacy of adversarial accuracy as the robustness metric for DL models, demonstrating that it fails to capture the full scope of a model’s resilience against attacks [11, 17, 18]. Robey et al. [17] have proposed to use the probability of misclassification as a more comprehensive measure of robustness, calculated through random sampling, although they acknowledge the computational challenges posed by the complexity of high-dimensional spaces. Oliver et al. [18] have explored non-adversarial spaces using geometric methods and assessed the extent of these regions within the sample space as an indicator of robustness. Despite these contributions, one can argue that neither misclassification probability nor adversarial region size sufficiently represent a model’s robustness, as they fail to consider the variation in a model’s response to varying levels of perturbation intensities.
@@
To overcome the shortcomings of existing robustness metrics, which focus solely on model robustness at a fixed perturbation level, we introduce a new metric called adversarial hypervolume. This metric evaluates models’ robustness by examining the normalized confidence scores of worst-case adversarial examples across various levels of perturbation intensities. It facilitates meaningful comparisons between defensive models with almost same adversarial accuracy and evaluation of the incremental robustness improvements provided by weaker defensive measures like Feature Squeezing and JPEG compression [19]. A higher adversarial hypervolume metric value indicates higher model robustness throughout various perturbation intensities, thereby offering a more comprehensive assessment of a model’s resilience against adversarial attacks. We illustrate a conceptual comparison between our proposed metric and two existing metrics in Figure 1.
@@
Our proposed metric, adversarial hypervolume, is based on the concept of hypervolume [20] from the multi-objective optimization literature [21, 22, 23]. We begin by constructing a multi-objective optimization problem that mirrors the complexity of actual attack scenarios, with the dual aims of minimizing the perturbation’s magnitude and simultaneously reducing the model’s confidence in the original classification. In contrast, much of existing research approaches the problem from the perspective of single-objective optimization with a distance constraint [7, 1] or employs a weighted sum of multiple objectives [11]. We subsequently delineate a set of adversarial examples that illustrate the model’s unique vulnerability landscape when subjected to an attack. Each example signifies a nadir of score corresponding to a particular magnitude of perturbation, thereby delineating the adversarial frontier. Finally, we quantify a model’s robustness by calculating the adversarial hypervolume, which measures the size of the region enclosed by the adversarial frontier.
@@
  • We introduce a new metric, the adversarial hypervolume, which quantifies the robustness of DL models across a range of perturbation intensities, thereby offering a comprehensive evaluation of a model’s resilience to adversarial attacks.
  • Utilizing this metric, we devise a straightforward method for its computation and implement an efficient training algorithm to improve the robustness of DL models.
  • Our empirical findings reveal that the adversarial hypervolume effectively measures the enhancement of robustness from weaker defensive mechanisms (i.e., input transformations) and delivers important supplementary insights, such as average confidence variation, which augment the understanding provided by adversarial accuracy.
  • We conduct extensive experiments to assess the robustness of cutting-edge state-of-the-art (SOTA) defensive models, establishing a benchmark for subsequent research.
@@ II-A Adversarial Attacks and Defenses
Adversarial Attacks. Following the identification of adversarial vulnerabilities in DL models [6], there has been a significant trend in research dedicated to the creation of adversarial attack algorithms. These studies explored to generate adversarial examples within different threat models [24, 25, 26]. Specifically, with different levels of information of the models, researchers have developed white-box attacks and black-box attacks respectively.
White-box Attacks. White-box attacks formulates an optimization problem aimed at creating adversarial examples by leveraging full model knowledge, e.g., gradient information. This approach has given rise to various attack methodologies, notably the PGD attack [1], the Carlini-Wagner (CW) attack [11], the FGSM attack [7], the Basic Iterative Method (BIM) attack [8], and the DeepFool attack [13].
Black-box Attacks. Black-box attacks operate under a threat model where only the model’s input and output are known [27, 28, 29, 30, 31, 32]. These attacks circumvent the limitations of white-box attacks in realistic scenarios where the model’s internal details are inaccessible. Notably, the Square attack [15], despite relying on less information than white-box attacks, proves to be more effective than some white-box methods. It has been incorporated into the AutoAttack framework [14] for robustness evaluation.
Attacks for Evaluating Robustness. Employing strong attacks to measure adversarial accuracies is a standard practice in robustness evaluation of DL models [1, 14]. Notably, the PGD attack [1] and the AutoAttack framework [14] stand as the most widely adopted attacks for such evaluations. AutoAttack combines the APGD attack with the Square attack. Moreover, the collection of the attack strategies is rapidly growing, now encompassing state-of-the-art approaches such as the Composite Adversarial Attack (CAA) [33] and AutoAE [34].
@@
Adversarial Defenses. Defensive strategies against adversarial attacks are designed to enhance a model’s inherent robustness or alleviate the effects of adversarial attacks. To improve the models’ robustness, a range of adversarial training techniques have been introduced [1, 35], along with fine-tuning on dataset or models [36, 37, 38]. Furthermore, several tactics have been devised to detect or neutralize adversarial examples, including input transformation [39, 19], adversarial purification [40, 41], and stateful defenses [42, 43]. Given that adversarial purification and stateful defenses are tailored for real-world classification systems rather than a standalone model, our study concentrates primarily on the robustness improvement brought by adversarial training and input transformation techniques.
@@
Adversarial Training. Adversarial training is a widely adopted method to enhance the robustness of DL models by integrating adversarial examples into the training dataset. This integration facilitates the model in developing more robust decision boundaries. Adversarial examples are often generated using the PGD attack [1], a substantial body of research supports its efficacy [44, 45, 46].
@@
Input Transformation. Input transformation is a cost-effective defensive strategy that can be seamlessly incorporated into systems with adversarially trained models to enhance robustness [39, 19]. However, the efficacy of input transformation has come into question, as it has been proven to be vulnerable to adaptive attacks [12]. Moreover, using adversarial accuracy to measure the robustness improvement from input transformation techniques may lead to misleading results, since the adversarial accuracy of disparate techniques tends to be low and indistinguishable. In contrast, our research introduces the adversarial hypervolume metric, which facilitates a comparative analysis of robustness improvements by providing more detailed insights into an averaged version of confidence value.
II-B Metrics for Evaluating Robustness
To gain a deeper insight into the robustness of models, researchers have introduced a variety of new metrics that extend beyond merely adversarial accuracy. Such metrics include the minimum adversarial perturbation [11], probabilistic robustness [17], and adversarial sparsity [18].
Minimal Perturbation. Quantifying the minimal magnitude of perturbation necessary to transform an input into an adversarial example is a pioneering attempt toward robustness evaluation [6]. Carlini et al. approached this issue by formulating an optimization problem that strikes a balance between perturbation magnitude and the confidence of the adversarial example, incorporating their linear combination [11]. However, this metric does not comprehensively reflect the full range of a model’s robustness across various perturbation levels.
Probabilistic Robustness. Robey et al. proposed a novel metric for evaluating probabilistic robustness, thereby redirecting attention away from traditional worst-case scenarios [17]. This metric assesses the probability of maintaining accurate predictions when a model is subjected to adversarial perturbations. However, one limitation is the considerable computational demand, typically requiring Monte Carlo simulations to estimate the distribution of adversarial examples, which becomes particularly challenging in high-dimensional settings.
@@ II-B Metrics for Evaluating Robustness
Adversarial Sparsity. Recent research by Oliver et al. has presented the concept of adversarial sparsity, which probes the boundary of non-adversarial regions to determine their proportion within the entire sample space [18]. Although the method aligns with the principles of probabilistic robustness, Oliver et al. suggest a geometric method to determine the adversarial sparsity, which yields improved convergence results.
II-C Multiobjective Optimization in Adversarial Robustness
Within the field of adversarial robustness, multiobjective optimization serves two main purposes: the enhancement of robustness and the development of a multiobjective adversarial framework. The former enhances model resilience against attacks by incorporating multiobjective optimization to refine the traditional norm-bounded problem. In contrast, the latter introduces new problem frameworks that more closely align with our research aims.
Multiobjecive Optimization for Robustness. Recent research efforts have advanced the application of multiobjective optimization in crafting adversarial examples. These efforts have expanded the original single-objective paradigm to include additional aims, yielding more diverse and robust adversarial examples [47, 48]. Additionally, multiobjective optimization has been employed in training models, offering a defensive enhancement to adversarial robustness [49].
Multiobjecive Adversarial Problem. The multiobjective adversarial problem introduces an innovative optimization framework to the domain of adversarial robustness, targeting the enhancement of model robustness across a range of perturbation intensities [50, 51, 52]. While Suzuki et al.’s [50] analysis aligns with our investigation, it restricts its scope to preliminary results with the VGG16 model. In contrast, Baia et al. [51] adopt non-norm-bounded attacks using established filters to generate adversarial examples, narrowing the broader applicability of their findings. Furthermore, Liu et al. [52] propose a novel approach for generating adversarial examples in NLP tasks with custom objectives which, however, suffers from limited applicability to the well-established classical norm-bounded attacks.
@@ III Adversarial Hypervolume
This section introduces the concept of adversarial hypervolume, a novel metric DL model robustness evaluation across varying perturbation intensities. Initially, we outline the multi-objective adversarial problem foundational to the adversarial hypervolume concept in Section III-A. Subsequently, we describe the adversarial frontier in Section III-B and detail the methodology for computing the adversarial hypervolume in Section III-C. We then explore the relationship between adversarial hypervolume, adversarial accuracy, and other established metrics in Section III-D.
@@ III-A Multiobjecive Adversarial Problem
Consider a classifier $f:\mathcal{X}\to\mathcal{Y}$, with $\mathcal{X}\subseteq[0,1]^{d}$ denoting the input space, and $\mathcal{Y}\subseteq[0,1]^{m}$ representing the output space, which signifies a probability distribution over $m$ class labels.
Typical Adversarial Problem. Adversarial attacks typically aim to determine the perturbation $\delta$, subjected to the constraint $\|\delta\|_{p}<\varepsilon$, where $\varepsilon$ is the perturbation budget, that can lead to a model’s misclassification. This has led to the formulation of the optimization problem below:
$\min_{\delta}\mathcal{L}_{\text{conf}}(f,x,y,\delta),\quad\text{s.t. }\|\delta\|_{p}\leq\varepsilon,$ (1)
where $\mathcal{L}_{\text{conf}}$ is the confidence loss and $p$ represents the norm used. As misclassification inherently involves a non-differentiable target, the confidence loss is utilized as a surrogate to estimate misclassification error. For the purposes of this paper, we adopt the minimization perspective, wherein lower $\mathcal{L}_{\text{conf}}$ indicate higher misclassification likelihood.
Multiobjecive Adversarial Problem. In this paper, we address a multiobjective adversarial problem that incorporates two objectives: the confidence loss and the distance loss. It is formulated as follows:
$\min_{\delta}\mathcal{L}(f,x,y,\delta)$ $=(\mathcal{L}_{\text{conf}}(f,x,y,\delta),\mathcal{L}_{\text{dist}}(\delta)),$ (2)
$\mathcal{L}_{\text{dist}}(\delta)$ $=\|\delta\|_{p}.$
This formulation delineates our goal to concurrently minimize these losses, which is more representative of real-world scenarios than the single-objective adversarial problem.
Marginal Confidence Loss. The confidence loss here, as mentioned above, is a surrogate of the misclassification error. We adopt the marginal confidence loss as a confidence proxy, defined as:
We have set the non-positive margin to be zero to omit the negative values, which are not indicative of misclassification. Other confidence loss functions can be used under our framework.